OpenStudy (chroncake):

3t + 2b = 9 1t + 2b=7 solve for t and b (:

KjSaif (kjsaif):

Are these two different equations? or both in one

OpenStudy (chroncake):

both in one

KjSaif (kjsaif):

ok

OpenStudy (rcp031):

Just convert any variable in terms of other n solve by method of substitution

OpenStudy (faiqraees):

multiply the second equation by -1 and then use elimination method

OpenStudy (rcp031):

Like t=7-2b

OpenStudy (chroncake):

so t=1 and b=3?

KjSaif (kjsaif):

Step 1: Solve2b+t=7for t: 2b+t+−2b=7+−2b(Add -2b to both sides) t=−2b+7 Step: Substitute−2b+72b+3t=9: 2b+3t=9 2b+3(−2b+7)=9 −4b+21=9(Simplify both sides of the equation) −4b+21+−21=9+−21(Add -21 to both sides) −4b=−12 −4b −4=−12/−4 (Divide both sides by -4)b=3 Step: Substitutet=−2b+7: t=−2b+7 t=(−2)(3)+7 t=1(Simplify both sides of the equation)

KjSaif (kjsaif):

b=3 and t=1

KjSaif (kjsaif):

You are correct :) @chroncake

OpenStudy (chroncake):

thanks guys! (:

OpenStudy (leenathan):

they are correct

OpenStudy (leenathan):

also welcome to open study